For equation, state the value of the discriminate and describe the nature of the solutions:
〖4x〗^2-6x+13=0
Discriminate =_________
2 real solutions or
1 real solution or
2 imaginary solutions

Answer by nyc_function(2741)

What is discriminate? I think you meant to type discriminant.
The discriminant is the term that lies in the radicand of the quadratic formula.
b^2 - 4ac = discriminant.
In this sample, a = 4, b = -6 and c = 13.
We now plug and chug.
(-6)^2 - 4(4)(13) =
36 - 4(52) =
36 - 208 = -172
Since the result is a negative number, there are two complex roots. In other words, there are 2 imaginary roots as solutions.